A solution of HClO4 was standardized by dissolving 0.4008 g of primary-standard-grade HgO in a solution of KBr:



HgO(s) + 4Br- + H2O  HgBr42- + 2OH-



The liberated OH- consumed 43.75 mL of the acid. Calculate the molar concentration of the HClO4.


1
Expert's answer
2021-11-24T12:21:19-0500

Solution:

Calculate the moles of HgO:

The molar mass of HgO is 216.59 g/mol

Hence,

(0.4008 g HgO) × (1 mol HgO / 216.59 g HgO) = 0.00185 mol HgO


Balanced chemical equations:

(1): HgO + 4Br+ H2O → HgBr42− + 2OH 

(2): OH + HClO4 → ClO4 + H2O


According to the chemical equation (1): 1 mol HgO produces 2 mol OH

Thus, 0.00185 mol HgO produces:

(0.00185 mol HgO) × (2 mol OH / 1 mol HgO) = 0.0037 mol OH


According to the chemical equation (2): 1 mol OH reacts with 1 mol HClO4

Thus, 0.0037 mol OH reacts with:

(0.0037 mol OH) × (1 mol HClO4 / 1 mol OH) = 0.0037 mol HClO4


Molarity of HClO4 = Moles of HClO4 / Volume of HClO4

Molarity of HClO4 = (0.0037 mol) / (0.04375 L) = 0.08457 M = 0.0846 M

Molarity of HClO4 = 0.0846 M


Answer: The molar concentration of HClO4 is 0.0846 M